OpenStudy (anonymous):

How do you write an equation in slope-intercept form for (-3,-1); m=4

OpenStudy (lgbasallote):

use point slope form y + 1 = 4(x + 3) y + 1 = 4x + 12 y = 4x + 11

OpenStudy (precal):

y=mx+ b -1=4(-3)+b -1=-12+b add 12 to both sides 11=b y=4x+11